OpenStudy (anonymous):

Anyone have any idea how to expand this with partial fractions? s/(s+1)^2(s^2+w^2)

OpenStudy (anonymous):

\[\frac{ s }{ (s+1)^2(s^2+\omega^2) }\]

OpenStudy (anonymous):

\[\frac{ s }{ (s+1)^2(s^2+w^2) }=s\times \frac{ 1 }{ (s+1)^2(s^2+w^2) }\] \[=s\times \frac{ 1}{ (s+1)^2 }\times \frac{ 1 }{ s^2+w^2 }\] \[=s\times (s+1)^{-2}\times (s^2+w^2)^{-1}\]

OpenStudy (anonymous):

Is this what you need to do?

OpenStudy (abb0t):

no.

OpenStudy (anonymous):

hmm I got to \[\frac{ s }{ (s+1)^2(s^2+\omega^2) }= \frac{ A }{ (s+1)^2 }+\frac{ B }{ (s+1) }....\]

OpenStudy (anonymous):

I don't know how to expand s^2 + w^2?

OpenStudy (anonymous):

Ah okay...

OpenStudy (abb0t):

\[= \frac{ A }{ (s+1) }+\frac{ B }{ (s+1)^2 }+\frac{ Cx+D }{ (s^2+w^2) }\]

OpenStudy (anonymous):

so that I can solve for a b and c given that c will be the expansion for s^2+w^2

OpenStudy (anonymous):

Yep Okay.

OpenStudy (anonymous):

oh snap that's what I wanted. How'd you know that it was cx+d?

OpenStudy (abb0t):

all polynomials are written in the form Cx+D.

OpenStudy (anonymous):

so if I want to solve for c and d would I just multiply both sides by cx+d? and let s=-cx-d?

OpenStudy (abb0t):

No.

OpenStudy (abb0t):

You are solving for A, B, C, and D.

OpenStudy (anonymous):

I know how to solve for b and a

OpenStudy (abb0t):

hmm....I usually get rid of the denominators, group like terms and factor out...then I solve for each individual variable, or I make a system of equations.

OpenStudy (abb0t):

You can try your method though see if it works

OpenStudy (anonymous):

would you consider omega as a constant?

OpenStudy (anonymous):

Oh I like your method better actually

hartnn (hartnn):

the degree of numerator polynomial is always 1 less than the degree of denominator, in partial fractions. thats why when we have linear expression in denominator, we write constants A,B,C... in numerator, so, when we have quadratic expression in denominator, we'll write a linear expression in numerator, like Ax+B,Cx+D,...and so on...

OpenStudy (abb0t):

omega?! where did Ω come from? Lol

hartnn (hartnn):

to find C and D, comparing the co-efficients is the best method...

hartnn (hartnn):

w <----omega

hartnn (hartnn):

and yes, its constant

OpenStudy (abb0t):

oOoooooH. LOL thought it was a w. Lol. yeah constant.

OpenStudy (anonymous):

oh so multiply to get same denominator and get rid of it then group terms based on degrees to solve for variables. Yeah gotcha thanks for your help I shall try that

OpenStudy (anonymous):

thanks for all your help

OpenStudy (abb0t):

yeah.